How would I solve the compound inequality?
4x-1/6 less than or equal to -5 or 4x-1/6 is greater than or equal to 5
The solution of the compound inequality is {x x less than or equal to __ or x greater than or equal to ___}

Answer by stanbon(75887)

4x-1/6 <= -5 or 4x-1/6 >= 5
-----------------------------
24x - 1 <= -30
24x <= -29
x <= -29/24
---------------------------
OR
24x -1 >= 30
24x >= 31
x >= 31/24
=========================
Cheers,
Stan H.
------

The solution of the compound inequality is {x x less than or equal to __ or x greater than or equal to ___}